# Higher-dimensional magnetic Skyrmions

### v1

@@ 2.1.3 Axially symmetric DMI
The Néel-type DM term (2.31) already possesses $\operatorname{SO}(3)$ symmetry, whereas the general Bloch-type generalization (2.7) does not. Considering the $\Theta$ tensor (2.15), we can relax the spherical (i.e. $\operatorname{SO}(3)$) symmetry restriction to only an axial symmetry. Using the transformation (2.10), but with the $\operatorname{SO}(4)$ transformation, i.e.,
@@
with $\rho^{\prime}\in\operatorname{SO}(2)$, $\mathbf{1}_{2}$ being the $2\times 2$ unit matrix, and with the $\operatorname{SO}(2)\subset\operatorname{SO}(3)$ spatial rotation
@@
The two vectors proportional to $\gamma$ and $\eta$ are invariants under the combined $\operatorname{SO}(2)$ rotation in the $x$-$y$ plane, but $(F^{1},F^{2})$ still transforms, so we must set them to zero.
@@ 2.2 Spherically symmetric magnetic Skyrmions
In order to construct a spherically symmetric magnetic Skyrmion, we necessarily need to consider only the $\operatorname{SO}(3)$ invariant subtensors, which for the Bloch-type DM term corresponds to the $\alpha$ and $\beta$ parts of Eq. (2.18). Considering the model that closest resembles the 2D magnetic Skyrmion, we take the energy density to be
@@
The kinetic term, ${\cal E}_{2}$, is invariant under $\operatorname{O}(4)_{\rm internal}\times\operatorname{O}(3)_{\rm space}\ltimes T_{3}$ symmetry, which is broken down to $\operatorname{O}(3)_{\rm internal}\times\operatorname{O}(3)_{\rm space}\ltimes T_{3}$ by the potential, ${\cal E}_{0}$, where $T_{3}$ are spatial translations and $\ltimes$ is the semi-direct product. The DM term, ${\cal E}_{1}$, further breaks the symmetry which is clear from the derivation in Sec. 2.1.1, viz. the internal symmetry is locked with the spatial rotations, yielding the symmetry of the theory
@@
More interesting is the $\operatorname{SO}(3)$-invariant DM term (2.18), which under the hedgehog Ansatz reduces to
@@
If we instead consider the generalized Néel-type DM term (2.31), we obtain under the hedgehog Ansatz the exactly the $\alpha$-part of the Bloch-type DM (see Eq. (2.41)), as expected. The symmetry is enhanced when $\delta=0$ and in that case the above density reduces to
@@ 2.2.1 Full equations of motion
The full equation of motion reads
@@ 2.2.2 Numerical Skyrmion solutions
We will now compute a spectrum of numerical solutions for the 3D generalization of the Skyrmion with a higher-dimensional DM term that is $\operatorname{SO}(3)$-invariant: namely the $\alpha$-invariant part of Eq. (2.18). We utilize the rescaling of lengths, so that the coefficient of the higher-dimensional DMI is unity, but the energy is measured in units of inverse $\kappa$, see Eq. (2.49). We solve the equation of motion (2.47) with $\kappa$ set to unity (under the rescaled scheme) using gradient flow. The gradient flow method is suitable for the stable 3D Skyrmions and is easier to use than a shooting method due to the large sizes of the Skyrmions, when $m$ is small. We have cross-checked the solutions using the shooting algorithm and confirmed their accuracy.
@@
Figure 1: Profile functions for 3D magnetic Skyrmions with higher-dimensional $\operatorname{SO}(3)$-invariant DM term. Left panels (a,c,e) show the profile on a regular scale and the right panels (b,d,f) on a log scale. The $m_{\rm crit}$ solution is shown with a dashed line. The rows of panels correspond to the potential power parameter $p=1,\tfrac{3}{2},2$. No solutions exist for $m>m_{\rm crit}$ nor for $m=0$.
@@
Figure 2: Energy density and topological charge density for 3D magnetic Skyrmions with higher-dimensional $\operatorname{SO}(3)$-invariant DM term. Left panels (a,c,e) show the total energy densities weighted with $r$ (for visualization purposes) as black lines and the DM contribution as red lines. Right panels (b,d,f) show the topological charge density, weighted with $r^{2}$. The $m_{\rm crit}$ solution is shown with a dashed line. The rows of panels correspond to the potential power parameter $p=1,\tfrac{3}{2},2$. No solutions exist for $m>m_{\rm crit}$ nor for $m=0$.
@@ 2.3.1 Numerical sphaleron solutions
We solve the equation of motion (2.47) with $\kappa$ set to unity (under the rescaled scheme) using the shooting method. The reason for changing the algorithm from gradient flow to the shooting method, is that the magnetic sphaleron solutions are unstable, which makes the gradient flow method easily flow to the stable 3D magnetic Skyrmions instead. Linearizing the equation of motion (2.47), we find that at small distances $r\ll 1$, the profile function behaves like $\chi=\pi-cr+\mathcal{O}(r^{2})$. We use $c$ as the shooting parameter and require $\chi(r_{\rm max})=0$, with a suitably large $r_{\rm max}$. The typical values of the shooting parameter for the magnetic sphalerons are in the range $c\approx 2\div 6$.
@@
Figure 3: Profile functions for magnetic sphalerons with higher-dimensional $\operatorname{SO}(3)$-invariant DM term, for (a) $p=1$, (b) $p=\tfrac{3}{2}$ and (c) $p=2$. The $m=0$ solution is shown with a green line and the $m_{\rm crit}$ solution with a dashed line. No solutions exist for $m>m_{\rm crit}$.
@@
Figure 4: Energy density and topological charge density for magnetic sphalerons with higher-dimensional $\operatorname{SO}(3)$-invariant DM term. Left panels (a,c,e) show the total energy densities weighted with $r^{2}$ as black lines and the DM contribution as red lines. Right panels (b,d,f) show the topological charge density, weighted with $r^{2}$. The $m=0$ solution is shown with a green line (light-green for the DM term) and the $m_{\rm crit}$ solution is shown with a dashed line. The rows of panels correspond to the potential power parameter $p=1,\tfrac{3}{2},2$.
@@
Figure 5: (a) Sizes and (b) energies of magnetic sphalerons (bottom series) versus 3D magnetic Skyrmions (top series) with higher-dimensional $\operatorname{SO}(3)$-invariant DM term. (a) The size is measured simply by the radius where $\chi=\frac{\pi}{2}$. (b) $-E$ is shown for the magnetic Skyrmions, as they have negative energy (i.e. they are stable solitons). Black pluses correspond to $p=1$, red crosses to $p=\tfrac{3}{2}$ and blue plus-crosses to $p=2$. The critical masses, $m_{\rm crit}$, correspond to the largest sphalerons and the smallest Skyrmions.
@@ 2.4 A connection to the Hopfion
so that the theory is $E[\bm{n}]+E_{0}^{\rm def}[\bm{n}]$, with $E[\bm{n}]$ of Eq. (2.1). This potential is compatible with the vacuum $\bm{N}=(0,0,0,1)^{\rm T}$, but squeezes the unbroken 2-sphere. In the limit $m^{\prime}\to 0$ we return to our higher-dimensional model, whereas in the limit $m^{\prime}\to\infty$, the target space is effectively reduced from the 3-dimensional $S^{3}$ to the 2-dimensional $S^{2}$. Being too energetically expensive to deviate from $n^{3}=0$, the theory with the $\operatorname{SO}(3)$-invariant higher-dimensional DM term (2.18) reduces in the $m^{\prime}\to\infty$ limit to
$E^{\rm red}[\bm{u}]=\int\mathop{}\!\mathrm{d}^{3}x\;\left[\frac{1}{2}\partial_{i}\bm{u}\cdot\partial_{i}\bm{u}+\kappa\alpha\bm{u}\cdot\nabla\times\bm{u}-\kappa(\beta-\alpha)\epsilon^{ij}u^{i}\partial_{3}u^{j}-m^{2}(1-u^{3})\right],$ (2.75)
@@
$E^{\rm red}[\bm{u}]=\int\mathop{}\!\mathrm{d}^{3}x\;\left[\frac{1}{2}\partial_{i}\bm{u}\cdot\partial_{i}\bm{u}+\kappa\bm{u}\cdot\nabla\times\bm{u}-m^{2}(1-u^{3})\right],$ (2.76)
@@ 3 A magnetic-QCD hybrid Skyrme model
We will now briefly consider the hybrid model consisting of the higher-dimensional magnetic Skyrmion model of Sec. 2 and the standard Skyrme model that is often used in a high-energy context [10, 11]. We do this by amending the energy (2.1) with the so-called Skyrme term, $E_{4}=\int\mathop{}\!\mathrm{d}^{3}{\cal E}_{4}$, whose density is [10, 11]
@@ 3.1 Derrick phase diagram
We notice, that if $\frac{1}{\tilde{e}^{2}}$ is too large, both saddle points are negative and only one solution exists. Setting $\kappa=\sqrt{8/3}$, the two saddle point coalesce. In this case, the condition $y(\widetilde{R}_{-}^{\rm saddle})>0$ becomes
@@ 5 Conclusion
In this paper, we have generalized the Dzyaloshinskii-Moriya interaction for a 3D magnetization vector to a 4D magnetization vector – hence a higher-dimensional version of the DMI. Using the property that it is first order in derivatives and must be nonvanishing, we deduced that it must be contracted with a tensor that is antisymmetric in $\operatorname{O}(4)$-indices and bears a 3D spatial index. Considering furthermore the simplified case of $\operatorname{SO}(3)$ invariance, we are able to reduce the DM term to just two invariant structures, that we denote the $\alpha$ and $\beta$ parts. It turns out that the $\beta$-part is irrelevant for the spherically symmetric case and only the $\alpha$-part stabilizes the solitons. Due to the higher dimension of space, there are in fact two solitons with just the DMI and a potential: the Skyrmion and a sphaleron, where the latter is an unstable and smaller solution of the model. Upon reduction to the Hopfion model, the $\beta$-part nevertheless plays an important role. Finally, we consider including the Skyrme term, thus considering a magnetic-QCD hybrid Skyrme model that turns out to host a large and a small Skyrmion as well as a sphaleron – at least in a small region of parameter space of the model. The large Skyrmion is stable, the small Skyrmion is metastable and the sphaleron is unstable. The phase diagram is quite rich and there exist lines where the different types of solutions coalesce and in the remaining part of the parameter space of the model, there exists only one solution – the Skyrmion.
@@
We have focused on spherically symmetric solitons in this paper. Using the spherically symmetric hedgehog Ansatz, we found that $\beta$-part of the $\operatorname{SO}(3)$-invariant DM term (2.18) vanishes and so we effectively only used the $\alpha$-part of the $\operatorname{SO}(3)$-invariant DM term. Deformed solitons, however, may feel the presence of the $\beta$-part of the DM term, which in principle could lower the mass of the soliton. We leave the exploration of such a possibility for future work.
We also considered a connection to the Hopfion in chiral magnets, and although there exists a higher-dimensional DM term and a potential with a limit that connects to the Hopfion model, the interpolation is not expected to be smooth. That is, the topology changes, sending the masses of the nontrivial $\pi_{3}(S^{3})$ solitons to infinity and the nontrivial $\pi_{3}(S^{2})$ solitons appear first as metastable and then finally become stable at the end point of the limit. A smooth interpolation is thus desirable and will be left for future work.
@@ Acknowledgments
S. B. G. thanks the Outstanding Talent Program of Henan University and the Ministry of Education of Henan Province for partial support. The work of S. B. G. is supported by the National Natural Science Foundation of China (Grant No. 12071111) and by the Ministry of Science and Technology of China (Grant No. G2022026021L). The work of S. B. is supported by the INFN special research project grant “GAST” (Gauge and String Theories).

### v2

@@ 2.1.3 Axially symmetric DMI
The Néel-type DM term (2.31) already possesses $\operatorname{SO}(3)_{\rm diag}$ symmetry, whereas the general Bloch-type generalization (2.7) does not. Considering the $\Theta$ tensor (2.15), we can relax the spherical (i.e. $\operatorname{SO}(3)$) symmetry restriction to only an axial symmetry. Using the transformation (2.10), but with the $\operatorname{SO}(4)$ transformation, i.e.,
@@
with $\rho^{\prime}\in\operatorname{SO}(2)_{t}$, $\mathbf{1}_{2}$ being the $2\times 2$ unit matrix, and with the $\operatorname{SO}(2)_{s}\subset\operatorname{SO}(3)_{s}$ spatial rotation
@@
The two vectors proportional to $\gamma$ and $\eta$ are invariants under the combined $\operatorname{SO}(2)$ rotation in the $x$-$y$ plane, but $(F^{1},F^{2})$ still transform, so we must set them to zero.
@@ 2.2 Spherically symmetric magnetic Skyrmions
In order to construct a spherically symmetric magnetic Skyrmion, we necessarily need to consider only the $\operatorname{SO}(3)_{\rm diag}$ invariant subtensors, which for the Bloch-type DM term corresponds to the $\alpha$ and $\beta$ parts of Eq. (2.18). Considering the model that closest resembles the 2D magnetic Skyrmion, we take the energy density to be
@@
The kinetic term, ${\cal E}_{2}$, is invariant under $\operatorname{O}(4)_{t}\times\operatorname{O}(3)_{s}\ltimes T_{3}$ symmetry, which is broken down to $\operatorname{O}(3)_{t}\times\operatorname{O}(3)_{s}\ltimes T_{3}$ by the potential, ${\cal E}_{0}$, where $T_{3}$ are spatial translations and $\ltimes$ is the semi-direct product. The DM term, ${\cal E}_{1}$, further breaks the symmetry which is clear from the derivation in Sec. 2.1.1, viz. the internal symmetry is locked with the spatial rotations, yielding the symmetry of the theory
@@
More interesting is the $\operatorname{SO}(3)_{\rm diag}$-invariant DM term (2.18), which under the hedgehog Ansatz reduces to
@@
If we instead consider the generalized Néel-type DM term (2.31), we obtain under the hedgehog Ansatz exactly the $\alpha$-part of the Bloch-type DM (see Eq. (2.41)), as expected. The symmetry is enhanced when $\delta=0$ and in that case the above density reduces to
@@ 2.2.1 Full equations of motion
The full equations of motion read
@@ 2.2.2 Numerical Skyrmion solutions
We will now compute a spectrum of numerical solutions for the 3D generalization of the Skyrmion with a higher-dimensional DM term that is $\operatorname{SO}(3)_{\rm diag}$-invariant: namely the $\alpha$-invariant part of Eq. (2.18). We utilize the rescaling of lengths, so that the coefficient of the higher-dimensional DMI is unity, but the energy is measured in units of inverse $\kappa$, see Eq. (2.49). We solve the equation of motion (2.47) with $\kappa$ set to unity (under the rescaled scheme) using gradient flow. The gradient flow method is suitable for the stable 3D Skyrmions and is easier to use than a shooting method due to the large sizes of the Skyrmions, when $m$ is small. We have cross-checked the solutions using the shooting algorithm and confirmed their accuracy.
@@
Figure 1: Profile functions for 3D magnetic Skyrmions with higher-dimensional $\operatorname{SO}(3)_{\rm diag}$-invariant DM term. Left panels (a,c,e) show the profile on a regular scale and the right panels (b,d,f) on a log scale. The $m_{\rm crit}$ solution is shown with a dashed line. The rows of panels correspond to the potential power parameter $p=1,\tfrac{3}{2},2$. No solutions exist for $m>m_{\rm crit}$ nor for $m=0$.
@@
Figure 2: Energy density and topological charge density for 3D magnetic Skyrmions with higher-dimensional $\operatorname{SO}(3)_{\rm diag}$-invariant DM term. Left panels (a,c,e) show the total energy densities weighted with $r$ (for visualization purposes) as black lines and the DM contribution as red lines. Right panels (b,d,f) show the topological charge density, weighted with $r^{2}$. The $m_{\rm crit}$ solution is shown with a dashed line. The rows of panels correspond to the potential power parameter $p=1,\tfrac{3}{2},2$. No solutions exist for $m>m_{\rm crit}$ nor for $m=0$.
@@ 2.3.1 Numerical sphaleron solutions
We solve the equation of motion (2.47) with $\kappa$ set to unity (under the rescaled scheme) using the shooting method. The reason for changing the algorithm from gradient flow to the shooting method, is that the magnetic sphaleron solutions are unstable, which makes the gradient flow method easily flow to the stable 3D magnetic Skyrmions instead. Linearizing the equation of motion (2.47), we find that at small distances $r\ll 1$, the profile function behaves like $\chi=\pi-cr+\mathcal{O}(r^{2})$. We use $c$ as the shooting parameter and require $\chi(r_{\rm max})=0$, with a suitably large $r_{\rm max}$. The typical values of the shooting parameter for the magnetic sphalerons are in the range $c\approx[2,6]$.
@@
Figure 3: Profile functions for magnetic sphalerons with higher-dimensional $\operatorname{SO}(3)_{\rm diag}$-invariant DM term, for (a) $p=1$, (b) $p=\tfrac{3}{2}$ and (c) $p=2$. The $m=0$ solution is shown with a green line and the $m_{\rm crit}$ solution with a dashed line. No solutions exist for $m>m_{\rm crit}$.
@@
Figure 4: Energy density and topological charge density for magnetic sphalerons with higher-dimensional $\operatorname{SO}(3)_{\rm diag}$-invariant DM term. Left panels (a,c,e) show the total energy densities weighted with $r^{2}$ as black lines and the DM contribution as red lines. Right panels (b,d,f) show the topological charge density, weighted with $r^{2}$. The $m=0$ solution is shown with a green line (light-green for the DM term) and the $m_{\rm crit}$ solution is shown with a dashed line. The rows of panels correspond to the potential power parameter $p=1,\tfrac{3}{2},2$.
@@
Figure 5: (a) Sizes and (b) energies of magnetic sphalerons (bottom series) versus 3D magnetic Skyrmions (top series) with higher-dimensional $\operatorname{SO}(3)_{\rm diag}$-invariant DM term. (a) The size is measured simply by the radius where $\chi=\frac{\pi}{2}$. (b) $-E$ is shown for the magnetic Skyrmions, as they have negative energy (i.e. they are stable solitons). Black pluses correspond to $p=1$, red crosses to $p=\tfrac{3}{2}$ and blue plus-crosses to $p=2$. The critical masses, $m_{\rm crit}$, correspond to the largest sphalerons and the smallest Skyrmions.
@@ 2.4 A connection to the Hopfion
so that the theory is $E[\bm{n}]+E_{0}^{\rm def}[\bm{n}]$, with $E[\bm{n}]$ of Eq. (2.1). This potential is compatible with the vacuum $\bm{N}=(0,0,0,1)^{\rm T}$, but squeezes the unbroken 2-sphere. In the limit $m^{\prime}\to 0$ we return to our higher-dimensional model, whereas in the limit $m^{\prime}\to\infty$, the target space is effectively reduced from the 3-dimensional $S^{3}$ to the 2-dimensional $S^{2}$. Being too energetically expensive to deviate from $n^{3}=0$, the theory with the $\operatorname{SO}(3)_{\rm diag}$-invariant higher-dimensional DM term (2.18) reduces in the $m^{\prime}\to\infty$ limit to
$E^{\rm red}[\bm{u}]=\int\mathop{}\!\mathrm{d}^{3}x\;\left[\frac{1}{2}\partial_{i}\bm{u}\cdot\partial_{i}\bm{u}+\kappa\alpha\bm{u}\cdot\nabla\times\bm{u}-\kappa(\beta-\alpha)\epsilon^{ij}u^{i}\partial_{3}u^{j}+m^{2}(1-u^{3})\right],$ (2.75)
@@
$E^{\rm red}[\bm{u}]=\int\mathop{}\!\mathrm{d}^{3}x\;\left[\frac{1}{2}\partial_{i}\bm{u}\cdot\partial_{i}\bm{u}+\kappa\bm{u}\cdot\nabla\times\bm{u}+m^{2}(1-u^{3})\right],$ (2.76)
@@ 3 A magnetic-QCD hybrid Skyrme model
We will now briefly consider the hybrid model consisting of the higher-dimensional magnetic Skyrmion model of Sec. 2 and the standard Skyrme model that is often used in a high-energy context [10, 11]. We do this by amending the energy (2.1) with the so-called Skyrme term, $E_{4}=\int\mathop{}\!\mathrm{d}^{3}x\,{\cal E}_{4}$, whose density is [10, 11]
@@ 3.1 Derrick phase diagram
We notice, that if $\frac{1}{\tilde{e}^{2}}$ is too large, both saddle points are negative and only one solution exists. Setting $\kappa=\sqrt{8/3}$, the two saddle points coalesce. In this case, the condition $y(\widetilde{R}_{-}^{\rm saddle})>0$ becomes
@@ 5 Conclusion
In this paper, we have generalized the Dzyaloshinskii-Moriya interaction for a 3D magnetization vector to a 4D magnetization vector – hence a higher-dimensional version of the DMI. Using the property that it is first order in derivatives and must be nonvanishing, we deduced that it must be contracted with a tensor that is antisymmetric in $\operatorname{O}(4)$-indices and bears a 3D spatial index. Considering furthermore the simplified case of $\operatorname{SO}(3)_{\rm diag}$ invariance, we are able to reduce the DM term to just two invariant structures, that we denote the $\alpha$ and $\beta$ parts. It turns out that the $\beta$-part is irrelevant for the spherically symmetric case and only the $\alpha$-part stabilizes the solitons. Due to the higher dimension of space, there are in fact two solitons with just the DMI and a potential: the Skyrmion and a sphaleron, where the latter is an unstable and smaller solution of the model. Upon reduction to the Hopfion model, the $\beta$-part nevertheless plays an important role. Finally, we consider including the Skyrme term, thus considering a magnetic-QCD hybrid Skyrme model that turns out to host a large and a small Skyrmion as well as a sphaleron – at least in a small region of parameter space of the model. The large Skyrmion is stable, the small Skyrmion is metastable and the sphaleron is unstable. The phase diagram is quite rich and there exist lines where the different types of solutions coalesce and in the remaining part of the parameter space of the model, there exists only one solution – the Skyrmion.
@@
We have focused on spherically symmetric solitons in this paper. Using the spherically symmetric hedgehog Ansatz, we found that $\beta$-part of the $\operatorname{SO}(3)_{\rm diag}$-invariant DM term (2.18) vanishes and so we effectively only used the $\alpha$-part of the $\operatorname{SO}(3)_{\rm diag}$-invariant DM term. Deformed solitons, however, may feel the presence of the $\beta$-part of the DM term, which in principle could lower the mass of the soliton. We leave the exploration of such a possibility for future work.
We also considered a connection to the Hopfion in chiral magnets, and although there exists a higher-dimensional DM term and a potential with a limit that connects to the Hopfion model, the interpolation is not expected to be smooth. That is, the topology changes, sending the masses of the nontrivial $\pi_{3}(S^{3})$ solitons to infinity and the nontrivial $\pi_{3}(S^{2})$ solitons appear first as metastable states and then finally become stable at the end point of the limit. A smooth interpolation is thus desirable and will be left for future work.
@@ Acknowledgments
S. B. G. thanks the Outstanding Talent Program of Henan University for partial support. The work of S. B. G. is supported by the National Natural Science Foundation of China (Grant No. 12071111) and by the Ministry of Science and Technology of China (Grant No. G2022026021L). The work of S. B. is supported by the INFN special research project grant “GAST” (Gauge and String Theories).